I need to know the result of (x+3)^5 using Newton's Binomial formula.

Answer by ewatrrr(24785)

(a+b)^5 = a^5 + 5a^4b + 10a^3b^2 + 10a^2b^3 + 5ab^4 + b^5
(x+3)^5 = x^5 +15x^4 + 90x^3 + 270x^2 + 405x + 243